Circular No. 7783

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PULSARS IN M62 B. A. Jacoby and A. M. Chandler, California Institute of Technol-

ogy (Caltech); D. C. Backer, University of California, Berkeley; and S. B. Anderson and S. R. Kulkarni, Caltech, report the discovery of three new binary millisecond pulsars in the globular cluster M62 (NGC 6266): “The discovery was made with the recently commissioned 100-m Green Bank Telescope (GBT). M62 was observed at a frequency of 1.4 GHz for 4 hr on 2001 Aug. 16 with the GBT and the Berkeley-Caltech Pulsar Machine (BCPM), a flexible 2×96-channel digital filterbank. Confirmation observations were carried out during the week of Dec. 4. We searched for pulsars at the dispersion measure of the three previously known pulsars in the cluster

(D’Amico et al. 2001, http://xxx.lanl.gov/abs/astro-ph/?0105122).

PSR J1701–3006D has a spin period of 3.418 ms, an orbital period of 1.118 days, and a minimum companion mass of 0.12 M_{⊙}. PSR J1701–3006E has a 3.234-ms spin period and is orbited by a companion of at least 0.03 M_{⊙} every 0.16 day. PSR J1701–3006F has a 2.295-ms spin period, an orbital period of 0.2 day, and a minimum companion mass of 0.02 M_{⊙}. M62 is now the third globular cluster containing six or more known radio pulsars.”

SUPERNOVA 2001ic IN NGC 7503 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ic (IAUC 7770), obtained by P. Berlind on 2001 Dec. 14.08 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova about a week after maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 13 262 km/s for the host galaxy, the supernova expansion velocity is ∼ 12 500 km/s for Si II (rest 635.5 nm). The spectral-feature age (cf. IAUC 7774) of SN 2001ic was 7 ± 2 days past maximum light.

R. Chornock and A. V. Filippenko, University of California, Berkeley, report that inspection of uncalibrated CCD spectra (range 330–1000 nm) obtained on Dec. 23 UT with the Keck-I telescope reveals that SN 2001ic is of type Ia, about a month past maximum brightness.

SUPERNOVA 2000cl IN NGC 3318

Corrigendum. On IAUC 7782, line 1, for he read she

2002 January 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green